• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:57:26 on May 5,2008

    Entry number 237069

    This entry is a followup to: 237065

    Followups:

    keyword=Loop 1 (LH2) fan not well

    The loop 1 (LH2) fan is ill: it is set to be 12% or 29Hz, but only runs at ~20 Hz. It was still normal at 8 am yesterday (entry 236871),
    but around 10 am, it already started to be irregular. Due to that,
    the temperature and HP heater all went out of the normal range.
    With data taking going on and target is stable, we will leave it as it is for now. Will look at when there is beam down time.
    Target operator should keep an eye on it and page me if there is
    any further changes.